OpenStudy (anonymous):

(u/a)+(v/d)=1, (u/d)+(v/a)=1. find u and v

OpenStudy (anonymous):

can someone please explain how to do this?

OpenStudy (australopithecus):

u/a = 1/2 u/d = 1/2 v/a = 1/2 v/d = 1/2

OpenStudy (australopithecus):

set a = 2 d = 2 v = 1 u = 1

OpenStudy (australopithecus):

1/2 + 1/2 = 1

OpenStudy (australopithecus):

or 0.5 + 0.5 = 1

OpenStudy (anonymous):

(u/a)+(v/d)=1 (1) (u/d)+(v/a)=1 (2) sorry i was going to write a detailed solution, but i have to go soon and do not have time rearrange equation (1) to find u in terms of a,d,v substitute into equation (2) (where ever you see u, you put your expression you just got from (1) ) rearrange to find v sub back into (1) to find u

OpenStudy (anonymous):

nice @Australopithecus

OpenStudy (australopithecus):

Look at what is changing in the equation and what is staying the same :)

OpenStudy (australopithecus):

then apply logic

OpenStudy (australopithecus):

but yes you can apply system of equations